OpenStudy (anonymous):

finding the length of an arc so my arc is 63/360 radius is 7 so 3.14 x 14=43.96 and then im stuck the last step im not sure what is done

OpenStudy (anonymous):

43.96 is for the full circumference of the circle. What you have is only 63/360th of the circle, so the ratio of the length of the arc over the whole circumference will be the same as the ratio of the angle captured on the whole 360degree circle

OpenStudy (anonymous):

i think i need to simplfy the answer as well

OpenStudy (anonymous):

80/ 360 Arc length = (p · 8) *Remember the radius = 4 so diameter = 8 Arc length = (25.12) Arc length = 5.58

OpenStudy (anonymous):

Yep, that's exactly it